### v2

# TeleRAG: Efficient Retrieval-Augmented Generation Inference with Lookahead Retrieval

Abstract.
[…]
RAG inference, Acceleration, Prefetching.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: submissionid: 164
1. Introduction
(a) Typical pipeline stages of a RAG application.
(b) Illustration of TeleRAG’s lookahead retrieval mechanism and comparison to different RAG systems. TeleRAG prefetches relevant retrieval data from CPU to GPU, overlaps data transfer with the pre-retrieval stage, and accelerates retrieval with GPU-CPU cooperation.
Figure 1. Overview of RAG pipeline stages and comparison of TeleRAG to baseline systems.
Retrieval-augmented generation (RAG) has emerged as a powerful technique to enhance large language models (LLMs) by integrating them with external databases (Gao et al., 2023b; Asai et al., 2024c; Lewis et al., 2020). During inference, RAG retrieves relevant content from external data sources, usually indexed as vector datastores, to mitigate issues such as hallucinations (Mallen et al., 2023; Ram et al., 2023; Khandelwal et al., 2019) and incorporate up-to-date or private information (Izacard et al., 2023; Min et al., 2023).
Modern RAG applications share two key characteristics. (1) RAG applications are built as modular pipelines as shown in Figure 1(a), where a single query undergoes multiple rounds of LLM calls and retrievals, each of which serves different functions to improve overall output quality. For example, query transformation rounds (Gao et al., 2023a; Zheng et al., 2023; Peng et al., 2023; Asai et al., 2024a) generally occur before retrieval to refine the user’s query with LLMs (pre-retrieval generation). (2) RAG’s datastores are typically large, supported by recent works demonstrating that larger datastore consistently improves the performance of RAG applications (Min et al., 2023; Khandelwal et al., 2019; Borgeaud et al., 2022; Hardt and Sun, 2023; Shao et al., [n. d.]).
These characteristics create several significant challenges for efficient RAG inference, especially in latency-sensitive applications such as customer chatbots (Chemmagate, 2024; Akkiraju et al., 2024; Tung, 2024), financial analysis (Malec, 2024; MyScale, 2024), and emergency medical diagnosis (Klang et al., 2024; GM and Dhar, 2024). First, the latency of RAG systems is inherently increased due to multiple rounds of LLM generation and retrieval. Second, although GPUs can speed up both LLM generation and retrieval, the combination of these processes poses significant memory requirements, often exceeding available GPU resources (see §3.1). Hence, it is expensive or even infeasible to fully host both operations on GPUs.
Consequently, in local or custom deployments, which are common for RAG applications with private or sensitive data, the retrieval datastore is often offloaded to CPU memory to alleviate GPU memory constraints. However, while CPU offloading resolves memory limitations, it significantly increases retrieval latency (see §3.1), diminishing overall system efficiency. This memory pressure is not limited to local deployments; even for serving scenarios in data centers, where many concurrent requests are processed together and GPUs are plentiful, managing large index sizes still presents a significant challenge for achieving high throughput with latency service level objectives (SLOs). Allocating additional GPU memory to the retrieval data store directly reduces the memory available for the KV cache of the LLM serving system, limiting the effective batch sizes (Kwon et al., 2023). In contrast, CPU retrieval would increase the per-request latency by a lot, highlighting the need for techniques that reduce GPU memory consumption while preserving low latency.
To address the latency challenge, several recent works have been proposed to accelerate RAG inference. These approaches include hardware accelerators for retrieval (Jiang et al., 2023d; Quinn et al., 2025), reuse of KV-cache (Jin et al., 2024a; Lu et al., 2024; Yao et al., 2024), and speculative document retrieval (Wang et al., 2024b; Zhang et al., 2024b). However, these works do not address the substantial memory demands associated with retrieval. Therefore, designing a memory-efficient system for low-latency RAG deployments is crucial.
In this paper, we argue that a promising direction to reduce GPU memory usage while maintaining low latency lies in improving the design of the retrieval index. Specifically, the inverted file index (IVF) (Sivic and Zisserman, 2003) reduces retrieval latency by pre-clustering data and limiting runtime search to relevant clusters. It offers a way to leverage GPU-accelerated search while not increasing memory consumption by dynamically transferring only the necessary clusters from the CPU to the GPU. However, this data transfer overhead is a key bottleneck, and needs to be minimized (Khan et al., 2024) (see §3.2).
Our proposal. To tackle latency and memory bottlenecks in RAG inference, we introduce a novel mechanism called lookahead retrieval, which predicts the subsets of IVF clusters that will likely be accessed during runtime and proactively transfers them from the CPU to the GPU before the retrieval stage.
The key observation behind lookahead retrieval is that there is substantial semantic overlap between queries before and after the pre-retrieval stage. Namely, IVF clusters used by the initial input query, which is available well before the retrieval stage, are also likely to be relevant to the actual retrieval query (see §3.3).
Using this insight, we propose TeleRAG, an efficient inference system designed to optimize RAG latency while minimizing GPU memory consumption. TeleRAG employs lookahead retrieval to preemptively load relevant IVF clusters onto the GPU, effectively hiding CPU-GPU data transfer overhead during concurrent LLM generation. As illustrated in Figure 1(b), this approach significantly reduces retrieval latency without exceeding GPU memory constraints, enabling efficient execution of RAG applications. To maintain retrieval accuracy while adopting lookahead retrieval, TeleRAG simultaneously searches for clusters that were not prefetched (i.e., mispredicted) from the CPU and merges the results. TeleRAG also supports batch and multi-GPU inference.
A key challenge in deploying lookahead retrieval is determining the optimal number of IVF clusters to prefetch: Prefetching too many clusters increases data-transfer overhead, whereas prefetching too few could result in higher retrieval latency due to increased CPU processing. To address this, we propose a profile-guided approach coupled with an analytical model that dynamically determines the ideal prefetch amount based on pipeline characteristics and hardware configurations.
[…]
In summary, we make the following key contributions:
  • Analyzing the correlation of the queries between the pre-retrieval generation and retrieval stages, revealing significant overlap in their corresponding IVF clusters.
  • Proposing lookahead retrieval, which prefetches likely IVF clusters to the GPU, and hides CPU-GPU data transfer time during pre-retrieval generation.
  • Developing TeleRAG, an efficient RAG inference system that integrates lookahead retrieval, resulting in significant acceleration of RAG with minimal GPU memory usage.
2. Background
Figure 2. Overview of RAG.
2.1. RAG
RAG is a technique that enhances the capabilities of LLMs by integrating them with information retrieval to generate more accurate and relevant text (Gao et al., 2023b; Asai et al., 2024c; Lewis et al., 2020). The core idea behind RAG is to augment the LLM with relevant information retrieved from a large corpus of documents, improving the LLM’s ability to answer questions without hallucinations (Mallen et al., 2023; Ram et al., 2023; Khandelwal et al., 2019) and generate contents based on up-to-date or private information (Izacard et al., 2023; Min et al., 2023).
RAG workflow. A basic RAG workflow involves several phases, including data store building, retrieval, and interactions with LLMs (Dixit, 2024; Ilin, 2023; Databricks, 2024; Wei, 2024; Vijeaswaran et al., 2024; Pathway, 2024; Gao, 2024). In order to build a data store, raw data in various formats is cleaned, converted to plain text, and divided into chunks. These chunks are then encoded into vectors using an embedding model and stored in a vector index, enabling efficient searching based on similarity. When a user provides a query, it is converted into a vector using the same embedding model, and the most similar chunks from the indexed database are retrieved. For a large-scale index, approximate algorithms such as the inverted file index (IVF) (Sivic and Zisserman, 2003) (see §2.2 for detail) are commonly used to accelerate the search process. The retrieved chunks, along with the original user query, are combined and given as a prompt to the LLM. The LLM then generates a response that relies on the information provided in the retrieved documents. This approach enables dynamic and informative responses in both single- and multi-turn dialogues (Lewis et al., 2020).
Modularity in modern RAG applications. However, naively integrating document retrieval into LLM generation can cause several issues. For example, the retrieval often struggles to find relevant content and may select irrelevant or mismatched chunks (Gao et al., 2023a; Zheng et al., 2023; Ma et al., 2023), and a single retrieval may not provide sufficient context, necessitating multiple retrieval rounds (Gao et al., 2023b).
To solve these issues, most state-of-the-art RAG models adopt a modular approach that employs multiple rounds of LLM calls and retrievals for a single query  (Gao et al., 2023b; Fan et al., 2024; Ilin, 2023; Gao, 2024; Databricks, 2024; Wei, 2024). Typically, they have the following types of steps (shown in Figure 2):
  • Pre-retrieval generation, used to assess whether retrieval is needed or to generate queries for retrieval. An example of a pre-retrieval technique is query transformation (Ma et al., 2023; Jagerman et al., 2023; Gao et al., 2023a; Zhou et al., 2022; Ye et al., 2023; Press et al., 2023; Peng et al., 2024; Zheng et al., 2023; Liu, 2022), which reformulates the original query to make it clearer and more suitable for the retrieval task.
  • Retrieval, used to identify relevant documents from the vector data store. This stage takes the output of pre-retrieval generation and generates data for the next stage.
  • Post-retrieval generation, generates response based on user query and retrieved documents. It can also perform additional process such as summarization (Jiang et al., 2023b; Kim et al., 2023b) or reranking (Zhuang et al., 2023; Stewart and Linsdell, 2024) on the retrieved documents.
  • Judgment, dynamically determines the execution flow. For example, it decides if more iteration is needed to enhance the response. Heuristics or LLMs can be used for judgement stage.
By proceeding through these functions, RAG applications can deliver more precise and contextually appropriate responses, significantly enhancing the capabilities of LLMs for various applications (Jin et al., 2024b; Gao et al., 2023b).
2.2. Vector Index and Inverted File (IVF)
The vector index is a crucial component of RAG applications that retrieves similar items from large datasets of high-dimensional vectors. Given the query vector $x\in\mathrm{R}^{D}$ and vector database $Y=\{y_{0},\ldots,y_{N-1}\}\subset\mathrm{R}^{D}$, which comprises $N$ vectors, the vector index aims to find the $k$ nearest neighbors of $x$ from the database:
$\text{k-argmin}_{i=0:N}d(x,y_{i}),$
where $D$ is the dimensionality of the vector determined by the embedding model, and $d$ denotes the distance function, which is typically the L2-norm or inner product (Johnson et al., 2019).
To improve search efficiency, the inverted file index (IVF) algorithm is widely used for large-scale vector databases due to its simplicity and effectiveness. IVF partitions the data store into clusters and restricts searches to the most relevant clusters, reducing the computational cost. To obtain the cluster assignment of each stored vector, it performs clustering algorithm such as $k$-means (Norouzi and Fleet, 2013) and partitions the database into $N_{c}$ clusters:
$\{C_{1},C_{2},\ldots,C_{N_{c}}\}=k\text{-means}(Y,N_{c}),$
where $C_{j}$ is the set of vectors assigned to the $j$-th cluster, and the cluster centroids $\{c_{1},c_{2},\ldots,c_{N_{c}}\}$ are obtained by taking the mean of each vectors in $\{C_{1},C_{2},\ldots,C_{N_{c}}\}$. Then, each database vector $y_{i}\in Y$ is assigned to the nearest cluster center:
$\text{Cluster}(y_{i})=\text{argmin}_{j=1:N_{c}}d(y_{i},c_{j}).$
With the trained centroids and cluster assignment, we can perform a more efficient index search. There are two steps involved in the IVF index searching. First, it will identify the nearest $L$ cluster centroids of a query vector $x$:
$\{c_{j_{1}},c_{j_{2}},\ldots,c_{j_{L}}\}=L\text{-argmin}_{j=1:C}\ d(x,c_{j}).$
This step is also referred to as the coarse-grained search in IVF, as it identifies candidates at the cluster level. Second, the fine-grained search is performed in the $L$ nearest clusters and identify $k$ closest vectors of $x$:
$k\text{-argmin}_{y_{i}\in\cup_{l=1}^{L}C_{j_{l}}}d(x,y_{i}).$
which involves sorting. In this way, IVF reduces search space and accelerates the retrieval process. Here, the hyperparameter $L$ from the first step is also referred as nprobe (Douze et al., 2024). When nprobe is larger, the IVF retrieval will be more accurate as it search more clusters. However, the retrieve latency is longer as more computation and data are needed.
Since the search process is highly parallelizable among each cluster and each vector, this search algorithm can be highly accelerated by GPUs. Open-source libraries offer efficient GPU implementations (Johnson et al., 2019; Rapidsai, 2022). However, IVF does not reduce the memory requirement for the index since the data for all clusters must be stored. As a result, IVF still requires a substantial GPU memory footprint, which becomes a bottleneck when GPU memory capacity is the constraint of RAG systems. Moreover, recent work reported that increasing the size of the data store positively affects RAG application performance (Min et al., 2023; Khandelwal et al., 2019; Borgeaud et al., 2022; Hardt and Sun, 2023; Shao et al., [n. d.]), exacerbating this issue.
2.3. Use Cases and Motivation
In this paper, we aim to achieve low-latency RAG inference without driving up GPU memory consumption, which is relevant for both local and data center settings. This goal is critical for efficiently serving a wide range of real-world applications where users expect near-instantaneous responses, thus imposing strict latency SLOs. Examples include customer chatbots (Chemmagate, 2024; Akkiraju et al., 2024; Tung, 2024; AWS, [n. d.]), financial analysis (Malec, 2024; MyScale, 2024), autonomous driving (Cai et al., 2024; Yuan et al., 2024), and emergency medical diagnosis (Klang et al., 2024; GM and Dhar, 2024).
On the one hand, there is significant demand for RAG applications that involve user-specific private data in local settings (Wang and Chau, 2024; Lam et al., 2023; Chung et al., 2023; Wutschitz et al., 2023; Ghodratnama and Zakershahrak, 2023). In such scenarios, the available GPU resources are usually limited, and the system serves a single request at a time. Therefore, CPU offloading is necessary when using a large vector index, which introduces substantial latency as we will show in the next section.
On the other hand, for large-scale RAG services in data centers (per, [n. d.]; gen, [n. d.]; Asai et al., 2024a), the goal is to serve many concurrent user requests cost-effectively, necessitating high system throughput. However, maximizing LLM serving throughput is often bottlenecked by the GPU memory available for the KV cache, which dictates the maximum possible batch size (Kwon et al., 2023). If a large retrieval datastore is loaded onto GPUs, it consumes memory that could otherwise be used for a larger KV cache, thereby limiting throughput. Hence, falling back to slow CPU retrieval to save GPU memory is a common option (Jiang et al., 2025), but this is not desirable for latency-sensitive applications due to latency overhead.
In both cases, a straightforward way to reduce inference latency is scaling up GPU resources to allow both the LLM and datastore to reside in GPU memory, which is usually cost-prohibitive. Therefore, in this paper, we aim to design an efficient RAG inference system that can satisfy latency SLOs without requiring a significant amount of GPU memory.
3. Analyzing Latency of RAG Pipelines
In this section, we analyze state-of-the-art RAG applications and identify their underlying systems challenges in achieving low latency. To conduct the analysis, we construct a 61 GB vector index with the Faiss library (Douze et al., 2024), and set the IVF clusters to 4096 following common practice (Asai et al., 2023). We use Llama-3-8B (AI@Meta, 2024) as the LLM, and run our analysis on an RTX 4090 (24 GB memory) and H100 (80 GB memory). §5.2 provides more details about our experimental setups.
3.1. End-to-end Latency of RAG Pipelines
(a) Batch = 1.
(b) Batch = 4.
Figure 3. Latency breakdown of six RAG pipelines on NQ dataset (Kwiatkowski et al., 2019) with one H100 GPU. nprobe is 256.
Figure 4. The breakdown of memory consumption at GPU and CPU for two different strategies: CPU offloading and GPU-based retrieval. The dotted line indicates the memory capacity of a RTX4090 GPU, which is a common used GPU for local deployment.
We first analyze the end-to-end latency of six representative RAG pipelines (see §5.1 for details) in two scenarios: (1) the LLM runs on the GPU, while retrieval is offloaded to the CPU, minimizing GPU memory usage, and (2) both the LLM and the vector index reside in GPU memory, enabling GPU-based retrieval for lower latency. If the GPU’s memory is insufficient, CPU offloading is necessary.
We set the nprobe to 256, a commonly used setting under this index’s configuration (see §5.2 for details), to measure retrieval latency. Figure 3 shows the breakdown of end-to-end latency into LLM inference part and retrieval part, based on 1024 randomly sampled queries from the NQ dataset (Kwiatkowski et al., 2019). We observe that, in the CPU-based retrieval baseline (first scenario), the retrieval phase dominates the latency, consuming 41.1% and 60.5% of total latency for batch sizes 1 and 4, respectively. In contrast, GPU-accelerated retrieval in the second scenario dramatically reduces this bottleneck, accounting for only 10.5% and 28.3% of the latency in comparable configurations. On average, GPU retrieval is $5.96\times$ (batch size 1) or $3.87\times$ (batch size 4) faster than CPU retrieval, reducing overall latency by approximately $1.5\times$ or $1.8\times$, respectively. Thus, accelerating retrieval on the GPU is crucial for improving end-to-end latency.
However, GPU acceleration comes with a significant memory cost. Figure 4 shows the memory requirements for both GPU and CPU. As Figure 4 shows, putting both LLM weights and the retrieval index on the GPU requires around 77 GB of memory, which exceeds the capacity of consumer GPUs like the RTX 4090 with 24 GB. Thus, GPU acceleration on retrieval is often not feasible when running on a lower-end GPU or indexing with a large datastore.
Even on GPUs capable of holding the entire index in memory (e.g., H100), the index’s memory footprint limits throughput in data center serving scenario. Batched inference maximizes throughput, but it is bottlenecked by the KV cache capacity for LLM serving (Kwon et al., 2023). For the case of the example above, storing the index on the GPU leaves minimal memory for the KV cache (e.g., ~3 GB, or ~20k tokens worth), restricting batch size, while offloading the index to the CPU significantly increases the available memory (e.g., to ¿500k tokens worth). This issue is exacerbated since RAG typically involves long contexts (Jin et al., 2024a; Lu et al., 2024; Yao et al., 2024) and there is a trend of growing model and index sizes (Min et al., 2023; Khandelwal et al., 2019; Borgeaud et al., 2022; Hardt and Sun, 2023; Shao et al., [n. d.]), increasing memory pressure and necessitating offloading. However, CPU offloading is suboptimal for latency-sensitive applications as discussed above.
Therefore, in the rest of this section, we try to answer the following question: Is it possible to achieve the latency of GPU-based retrieval while using much less GPU memory?
Figure 5. Latency breakdown of CPU-offload and runtime-fetch GPU retrieval, averaged over 512 random NQ queries.
3.2. Opportunities of GPU-accelerated Retrieval with Limited Memory
A straightforward approach to enable GPU retrieval with limited GPU memory is to fetch the necessary data from CPU to GPU on-demand for each query, leveraging the IVF index structure that narrows the search space (§2.2). While this method enables faster searching on the GPU, data fetching becomes the bottleneck.
In Figure 5, we compare the latency of the on-demand fetching system against CPU retrieval on an RTX 4090 GPU. We show three different nprobe values that determine the amount of data fetched. Overall, fetch time dominates latency due to the limited PCIe bandwidth between the CPU and GPU (32 GB/s). Although the GPU search is substantially faster than the CPU counterpart, the fetch overhead results in a higher end-to-end latency, which is about $3\%$ slower on average across nprobe values. Thus, to realize any meaningful speedup with this approach, the data fetching latency must be effectively hidden.
To hide data fetch costs, CPU-to-GPU transfer must precede retrieval, which requires predicting which data will be accessed. Fortunately, as we show next, the modern RAG pipelines offer a valuable hint as a query at the previous step.
Dataset HyDE SubQ Iter IRG FLARE S-RAG
(Gao et al., 2023a) (Liu, 2022) (Liu, 2022) (Shao et al., 2023) (Jiang et al., 2023c) (Asai et al., 2023)
NQ (Kwiatkowski et al., 2019) 73.1% 63.2% 91.5% 83.8% 79.1% 100.0%
HotpotQA (Yang et al., 2018) 75.3% 62.5% 89.6% 89.4% 80.2% 100.0%
TriviaQA (Joshi et al., 2017) 73.1% 61.6% 86.2% 86.1% 81.7% 100.0%
Table 1. IVF cluster overlapping rate between the input and output of the pre-retrieval generation. The nprobe is set to 512. Since Self-RAG does not incorporate query transform, its coverage is always $100\%$.
3.3. Overlapping of IVF Clusters
While the exact data to be retrieved is only known after the pre-retrieval generation is done, we observe that there are high similarities of the IVF clusters assignments between the queries at different stages.
Similarity of queries at different stages. During the pre-retrieval process of RAG pipelines (e.g., query transformation (Ma et al., 2023; Jagerman et al., 2023; Gao et al., 2023a; Zhou et al., 2022; Ye et al., 2023; Press et al., 2023; Peng et al., 2024; Zheng et al., 2023; Liu, 2022)), an LLM call is issued to refine an initial user query $q_{\text{in}}$ into a transformed query $q_{\text{out}}$, which is then used for retrieval. This process often rewrites the query into a different format (Gao et al., 2023a) or simplifies its complexities (Liu, 2022), which intuitively preserves the query’s core semantic content. Hence, the embedding vectors of $q_{\text{in}}$ and $q_{\text{out}}$ are likely to be similar. This similarity, in turn, suggests that the IVF clusters to which these queries would be assigned will overlap. Therefore, $q_{\text{in}}$ can serve as a valuable hint for predicting subsequent memory accesses.
Prediction coverage. To verify this hypothesis, we evaluate the average cluster coverage rate between prefetched clusters and clusters used for retrieval in three popular question-answer datasets (NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and TriviaQA (Joshi et al., 2017)) and six RAG pipelines. Table 1 shows the coverage when we prefetch $256$ clusters. From the table, we observe that IVF cluster overlap rates are consistently high across a range of datasets and pipelines. For instance, even the lowest reported values remain above 61.6% (for SubQ).
Opportunity. This data shows an opportunity for predicting required clusters, which can make it possible to hide data transfer overhead during LLM generation. In this paper, we aim to leverage this observation to accelerate the inference latency for RAG.
Figure 6. (a) The overview of lookahead retrieval compared against CPU-offloaded retrieval, and (b) the system design of TeleRAG. After identifying clusters ($C_{\text{in}}$) for the initial query $q_{\text{in}}$, $C_{\text{in}}$ is transferred to the GPU while concurrently generating $q_{\text{out}}$. At retrieval time, the GPU searches prefetched clusters while the CPU handles remaining ones, before results are merged.
4. Design of TeleRAG
Based on the high IVF cluster overlapping observation presented in §3.3, we propose TeleRAG, an efficient RAG inference system that utilizes lookahead retrieval to prefetch a minimal set of likely IVF clusters to GPU, and accelerate the retrieval process. In this section, we detail the design of lookahead retrieval and implementation of TeleRAG.
4.1. Overview
Figure 6 presents the overview of TeleRAG with lookahead retrieval. In Figure 6, $q_{\text{in}}$ denotes the initial input to the pre-retrieval stage and $q_{\text{out}}$ denotes its output, i.e., the query given to the subsequent retrieval stage. We highlight the IVF clusters corresponding to $q_{\text{in}}$ and $q_{\text{out}}$ with a red background ($C_{\text{in}}$), and a blue background ($C_{\text{out}}$), respectively. Due to the semantic similarity of $q_{\text{in}}$ and $q_{\text{out}}$, there is significant overlap between $C_{\text{in}}$ and $C_{\text{out}}$, marked with a purple background ($C_{\text{overlap}}$).
Given this, the proposed lookahead retrieval technique in TeleRAG operates in three key steps:
  1. (1)
    Pre-retrieval & data transfer: During LLM generation, identify $C_{\text{in}}$ and transfer corresponding data to GPU memory, prioritizing clusters closest to $q_{\text{in}}$. This leverages GPU DMA for concurrent transfer of clusters data and compute of LLM generation.
  2. (2)
    GPU similarity search: After identifying $C_{\text{out}}$, the GPU efficiently performs a similarity search on the predicted clusters ($C_{\text{overlap}}$) already in its memory.
  3. (3)
    CPU similarity search: Concurrently with the GPU search, the CPU processes the similarity search for the missed clusters ($C_{\text{miss}}$) that were not prefetched to the GPU, reducing the CPU’s overall workload.
In summary, our lookahead retrieval speeds up retrieval by overlapping data prefetching to the GPU with LLM generation and performing concurrent similarity searches on both the GPU (for predicted clusters) and CPU (for the remainder), significantly reducing the CPU’s computational workload.
4.2. Finding the Optimal Prefetch Amount
A key challenge in TeleRAG is to balance the benefit of reducing retrieval latency by prefetching data against the overhead of CPU-GPU transfers. Prefetching more clusters reduces the subsequent retrieval time but can also extend the transfer phase; if it extends beyond the LLM generation window, we lose the advantage of overlap and potentially introduce additional delay. Still, additional delays in data transfer may be worthwhile if they substantially reduce retrieval latency. To guide the choice of the optimal amount of data to prefetch, we develop a mathematical model based on profile information.
Mathematical model. Here, we denote $b_{p}$ as the number of bytes to prefetch and $B$ as the CPU-GPU bandwidth. The optimal amount of data to prefetch is denoted as $b_{p}^{*}$. To start, we let $t_{1}$ represent the combined time of prefetching and pre-retrieval LLM generation, and $t_{2}$ represents the retrieval time. We have: $t_{1}=\max(t_{\text{LLM}},t_{p})$ and $t_{2}=\max(t_{c},t_{g})$, where $t_{\text{LLM}}$ is the time for LLM generation, $t_{p}$ is prefetching time, $t_{c}$ is the CPU retrieval time, and $t_{g}$ is the GPU retrieval time. The objective is to minimize $t_{1}+t_{2}$.
Since prefetching time $t_{p}$ is proportional to $b_{p}$, we express $t_{1}$ as a piecewise function:
(1) $t_{1}=\begin{cases}t_{\text{LLM}},&\text{if }b_{p}\leq B\cdot t_{\text{LLM}},\\[6.0pt] \frac{b_{p}}{B},&\text{if }b_{p}>B\cdot t_{\text{LLM}},\end{cases}$
From Eq. 1, if we prefetch fewer bytes than can be transferred during LLM generation, $t_{1}$ is effectively just $t_{\text{LLM}}$ because the transfer overlaps completely with LLM execution.
As shown in §3.1, GPU retrieval ($t_{g}$) is generally much faster than CPU retrieval ($t_{c}$). Thus, we assume $t_{c}\gg t_{g}$. As CPU has a limited parallelism, $t_{c}$ usually grows proportionally with the number of clusters to be processed (Jiang et al., 2023a):
(2) $t_{2}=t_{c}=r_{\mathrm{miss}}\times n_{\mathrm{probe}}\times t_{cc},$
where $r_{\mathrm{miss}}$ is the miss rate (percentage of IVF clusters are not caught on GPU), $n_{\mathrm{probe}}$ is the total number of clusters to search, and $t_{cc}$ is the CPU time to search a single cluster. Increasing $b_{p}$ can only decrease or maintain the current miss rate, i.e., $\frac{\mathrm{d}r_{\mathrm{miss}}}{\mathrm{d}b_{p}}\leq 0$. Moreover, because clusters are prefetched in order of descending likelihood, we assume $r_{\mathrm{miss}}$ is either linear or a concave up function [^1] An upward U-shaped function whose second derivative is positive. of $b_{p}$, i.e., $\frac{\mathrm{d}^{2}r_{\mathrm{miss}}}{\mathrm{d}b_{p}^{2}}\geq 0$.
We now examine two cases:
  • Case 1: $b_{p}^{*}\leq B\cdot t_{\text{LLM}}$. Here, $t_{1}=t_{\text{LLM}}$ is constant because prefetching is fully overlapped with LLM generation. Since increasing $b_{p}$ in this regime will not increase $t_{1}$ and cannot worsen the miss rate, pushing $b_{p}$ to the boundary $B\cdot t_{\text{LLM}}$ minimizes $t_{1}+t_{2}$. Hence,
    (3) $b_{p}^{*}=B\cdot t_{\text{LLM}}.$
  • Case 2: $b_{p}^{*}>B\cdot t_{\text{LLM}}$. In this region, $t_{1}$ grows linearly with $b_{p}$, and we have: $\frac{\mathrm{d}^{2}}{\mathrm{d}b_{p}^{2}}(t_{1}+t_{2})=\frac{\mathrm{d}^{2}r_{\mathrm{miss}}}{\mathrm{d}b_{p}^{2}}\cdot n_{\mathrm{probe}}\cdot t_{cc}\geq 0.$ Therefore, $t_{1}+t_{2}$ is concave up, allowing at most one minimum. At the minimum point, we have:
    (4) $\frac{\mathrm{d}}{\mathrm{d}b_{p}}(t_{1}+t_{2})=0\implies\frac{1}{B}+\frac{\mathrm{d}r_{\mathrm{miss}}}{\mathrm{d}b_{p}}\cdot n_{\mathrm{probe}}\cdot t_{cc}=0.$
    From this, we obtain:
    (5) $b_{p}^{*}=B\times n_{\mathrm{probe}}\times t_{cc}\times\Delta r_{\mathrm{miss}},$
    where $\Delta r_{\mathrm{miss}}$ is the decrement of the miss rate for this round. If $b_{p}^{*}$ is indeed larger than $B\cdot t_{\text{LLM}}$, it becomes the global minimum; otherwise, the solution reverts to Case 1.
In summary, our analysis shows that the optimal prefetch amount $b_{p}^{*}$ can only lie at one of two points: (1) Prefetch until LLM generation completes (i.e. $b_{p}=B\cdot t_{\text{LLM}}$). (2) A point determined by Eq. 5. However, under typical CPU-GPU bandwidth (e.g., 55 GB/s on PCIe 5), the time spent loading additional clusters often outweighs any retrieval latency reduction from lowering $r_{\mathrm{miss}}$. Consequently, the second scenario in Eq. 5 becomes nearly infeasible in practice. Therefore, on current hardware, prefetching exactly until LLM execution ends is generally the most effective choice.
Profiling-guided approach. Although $b_{p}^{*}$ can be derived from the analysis above, it depends on knowing $t_{\mathrm{LLM}}$ for each query, which cannot be obtained ahead of time. To address this, we use a profiling-guided approach, leveraging the observation that, despite differences in query content, the output length (and thus generation time) for a RAG pipeline often remains similar across most queries. Accordingly, for each RAG pipeline, we can measure $t_{\mathrm{LLM}}$ on a calibration set containing n queries, and get estimated $\hat{b}_{p}^{*}=B\cdot\overline{t}_{\mathrm{LLM}}$, where $\overline{t}_{\mathrm{LLM}}=\mathrm{mean}\{t_{\mathrm{LLM},1},t_{\mathrm{LLM},2},\dots,t_{\mathrm{LLM},n}\}.$ This estimated $\hat{b}_{p}^{*}$ can then be used for incoming queries, ensuring a near-optimal prefetch amount.
4.3. Design Details
Sorting Optimization. TeleRAG accelerates the final $k$-argmin sorting step of IVF search by leveraging GPU. Because each cluster may contain thousands of points, sorting is particularly well-suited to GPU parallelism (Arkhipov et al., 2017). To enable this, TeleRAG transfers the scalar distance values computed on the CPU for $C_{\text{miss}}$ to the GPU. Unlike full vector data, this transfer incurs negligible overhead. The GPU then performs a global sort over the combined distances from $C_{\text{overlap}}$ and $C_{\text{miss}}$, delivering end-to-end speedups than the CPU-only retrieval systems.
Prefetch Target. TeleRAG’s prefetching mechanism targets a specific number of bytes ($b_{p}$) rather than the number of clusters to ensure more predictable performance. Since IVF cluster sizes are often highly uneven, targeting a fixed cluster count can lead to unstable data loading times, whereas a byte limit provides a clearer upper bound on the transfer duration based on bandwidth.
When selecting clusters to meet the prefetch byte budget ($b_{p}$), the system adds whole clusters sequentially based on proximity to the query. If adding the next closest cluster would exceed the budget, it is omitted entirely rather than being partially loaded, ensuring a clean division where each cluster is processed exclusively by either GPU or CPU.
Multi-Round Support. For multi-round RAG involving the same input query, the system performs a full prefetch (up to the budget) only in the first round, leveraging high cluster similarity across rounds. In subsequent rounds, it incrementally fetches only the additional required clusters that were not loaded previously, optimizing data transfer while respecting the memory budget.
Batch Support. For batch inference, a challenge for lookahead retrieval is that each query requires a different set of IVF clusters. In our design, we fix the prefetching budget to the amount we got from §4.2 irrespective of batch size, and equally distribute the budget to each query in a batch. As a result, although the average cluster hit rate per request will decrease for larger batch sizes, we can ensure each query’s retrieval search time is equally accelerated, forming a balanced batched retrieval.
Multi-GPU support. For multi-GPU systems, we perform a greedy search based on embedding similarities of queries to group similar queries into mini-batches, and distribute the mini-batches among GPUs. The greedy search is efficient and incurs negligible overhead: In our profiling for up to the batch size of 256, the latency to perform greedy search is less than $0.1$ s, which is minimal compared to the end-to-end latency of each batch.
4.4. Implementation Details
TeleRAG is implemented in Python and also leverages PyTorch’s (Ansel et al., 2024) operator ecosystem for efficient computation. The datastore index is initially constructed using FAISS (Douze et al., 2024), and its data structures, such as IVF centroids and cluster data, are converted to PyTorch tensors.
At runtime, cluster data is loaded into a contiguous pinned memory region on the CPU, enabling non-blocking memory copies to the GPU. A fixed-size contiguous buffer on the GPU is allocated based on the user’s configuration or GPU memory capacity during runtime.
To enable concurrent CPU-GPU data transferring and LLM generation, we utilize the _copy(non_blocking=True) API in PyTorch and use separate CUDA streams for prefetching and LLM. In this way, the data copy operations will not block the GPU to perform computation, and thus the TeleRAG can do prefetching in parallel to the pre-retrieval LLM generation.
To implement the index search with GPU-CPU cooperation, for the GPU part, we use a single matrix-vector multiplication that computes distances for all prefetched vectors; for the CPU part, we utilize multithreading in Python to parallelize similarity searches across clusters. We then move the distances computed from CPU to GPU, merge with distances on GPU and perform global sorting on GPU.
Specification Value
Dataset wiki_dpr (Karpukhin et al., 2020)
Dataset size 2.1 billion tokens
#\# of chunks 21 million
#\# of IVF cluster 4096
Embed model Contriever (Izacard et al., 2021)
Embed dimension 768
Index type FLAT [^2] Original embedding without compression for the best retrieval precision.
Distance metric Inner Product
Index size 61 GB
Table 2. Detailed configurations of our retrieval index.
5. Evaluation
We conduct extensive experiments to evaluate the effectiveness of TeleRAG. In this section, we describe the necessary details on how we set up the evaluations, present experimental results, and provide in-depth analysis and discussions.
5.1. Evaluation Datasets and RAG Models
Datastore. We build a datastore based on the wiki_dpr dataset (Karpukhin et al., 2020), a popular dataset that contains 2.1 billion tokens from Wikipedia. Following previous works (Asai et al., 2023; Karpukhin et al., 2020; Min et al., 2023), we chunk the passages by every 100 tokens, and use Contriever (Izacard et al., 2021) to generate an embedding for each chunk. The embeddings have a hidden dimension of 768.
Vector index. For the baseline retrieval, we build an IVF vector index using Faiss (Douze et al., 2024) on the datastore. As described in §4.4, we convert the Faiss index to a customized index in PyTorch for TeleRAG. See Table 2 for the detailed configurations of our vector index and datastore.
LLMs. We evaluate TeleRAG on the Llama model family (Touvron et al., 2023) in three different sizes (Llama-3.2-3B, Llama-3-8B, Llama-2-13B) to represent different use cases.
Setup Desktop Server
CPU Threadripper 5975 EPYC 9554
CPU memory size 512 GB 1.5 TB
GPU RTX4090 (NVIDIA, 2024a) H100 (NVIDIA, 2024b)
GPU memory size 24 GB 80 GB
CPU-GPU Bus PCIe 4 PCIe 5
Bandwidth n 32 (24) GB/s n n 64 (51) GB/s n
Table 3. Hardware specifications for our setups. In bandwidth, the number in the parentheses is the actual bandwidth we measured from our system.
RAG pipelines. We evaluate TeleRAG with six popular RAG pipelines. Figure 7 shows the overview of each pipeline. Note that although some pipelines lack pre-retrieval generation, the post-retrieval generation serves similar functionality for the retrieval of the next iteration. Below are brief descriptions of the evaluated RAG pipelines.
  1. (1)
    HyDE (Gao et al., 2023a) prompts LLM to generate a hypothetical paragraph and perform retrieval based on the embedding of the generated paragraph.
  2. (2)
    SubQestion (SubQ) [^3] Implemented in LlamaIndex as SubQuestionQueryEngine. (Liu, 2022) prompts LLM to generate multiple sub-questions and performs retrievals for each generated sub-question.
  3. (3)
    Iterative (Iter) [^4] Implemented in LlamaIndex as MultiStepQueryEngine. (Liu, 2022) prompts LLM to generate narrower questions first and iteratively refine them based on previous answers. At the end of each iteration, it prompts LLM to judge if the answer is good enough.
  4. (4)
    Iter-RetGen (IRG) (Shao et al., 2023) iteratively do retrieval and LLM generation for 3 iterations.
  5. (5)
    FLARE (Jiang et al., 2023c) iteratively issues retrievals based on the confidence (probability score) of predicted tokens for the upcoming sentence.
  6. (6)
    Self-RAG (S-RAG) [^5] We evaluate on the short-form version which only has one iteration. (Asai et al., 2023) uses the LLM to judge for retrieval, generate responses, and self-critique on the responses. We use fine-tuned models based on Llama2-7B and Llama2-13B from their official repository (Asai et al., 2024b).
Figure 7. Overview of six RAG pipelines that we evaluate.
Evaluation Datasets. We use three commonly used question-answering datasets, NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and TriviaQA (Joshi et al., 2017). For each dataset, we randomly sample 1024 queries and report the average unless otherwise specified.
5.2. Experiment Setups
Hardware setups. We evaluate TeleRAG on two hardware environments, Desktop and Server, which are equipped to represent the settings for the desktop and data center use cases. The Desktop has the RTX4090 (24 GB memory), and we use 3B and 8B models. The Server has the H100 (80 GB memory), and we use 8B and 13B models. These sizes represent common model sizes for RAG applications (Asai et al., 2023, 2024a). We do not evaluate the 13B model on Desktop as its model size (26 GB) exceeds RTX4090’s memory capacity. Table 3 summarizes the hardware configurations.
Nprobe and top-$k$. For the IVF index, nprobe is a hyperparameter that controls the trade-off between search efficiency and retrieval quality. A common heuristic is to set nprobe to $4\sqrt{N_{c}}$ (Zilliz, 2020). Given our index size of $N_{c}=4096$, we use nprobe = 256 ($4\sqrt{4096}$) by default, unless otherwise specified. For retrieval, top-$k$ denotes the number of most relevant documents to return. We use top-$k$ = 3, which is a standard choice in RAG.
RAG pipeline implementation. We implement the RAG pipelines with the FlashRAG framework (Jin et al., 2024b). For IRG, FLARE, and S-RAG, we use the framework’s default implementations. For the other pipelines, we reimplement them using FlashRAG’s APIs.
Benchmark methodology. We follow a benchmarking methodology from SGLang (Zheng et al., 2024) and use GPT-3.5-Turbo (OpenAI, [n. d.]) to run through each pipeline and record the input and output text of each step. During latency evaluation, we set the LLM’s output length based on the recorded text. This way, we ensure a fair latency comparison across different LLM models.
(a) End-to-end latency speedup with Llama-3.2-3B.
(b) End-to-end latency speedup with Llama-3-8B.
Figure 8. End-to-end latency speedup of TeleRAG and baseline on six RAG pipelines and three datasets, with RTX4090 GPU. nprobe is 256.
Baseline systems. To evaluate the latency of each pipeline, we construct a clean execution flow in Python that only contains LLM generation, datastore retrieval, and other necessary logical operations to fulfill each pipeline. For LLM generation, we use SGLang (Zheng et al., 2024), which is a state-of-the-art LLM inference engine. For retrieval, we use a popular retrieval library, Faiss (Douze et al., 2024), as the CPU-offloaded baseline. [^6] This baseline resembles frameworks like LlamaIndex and LangChain, which primarily function as wrappers around LLM serving and retrieval libraries. We build a system from the ground up to avoid the overhead introduced by their complex layers designed for advanced functionality.
Prefetching budget setups. Based on the methodology we described in §4.2, we profile each RAG pipeline with a small amount of 64 random samples from NQ (Kwiatkowski et al., 2019) and derive the prefetching budget of each pipeline.
(a) End-to-end throughput with six pipelines and different batch sizes using Llama-3-8B.
(b) End-to-end throughput with six pipelines and different batch sizes using Llama-2-13B.
Figure 9. End-to-end throughput of TeleRAG and the baseline across six RAG pipelines on the NQ dataset using Llama-3-8B and Llama-2-13B at different batch sizes on a H100 GPU. The nprobe is set to 256, and the $x$-axis represents the batch size.
Max prefetching memory. We set a maximum GPU memory limit for prefetching in each configuration. For Server, we allocate up to 16 GB. For Desktop, we allocate up to 10 GB and 3.75 GB for the 3B and 8B models, respectively. These settings demonstrate that TeleRAG can efficiently operate using only a small fraction (up to 40% for RTX4090 and 20% for H100) of total GPU memory.
5.3. Evaluation Results
Single-query latency on RTX4090. We evaluate the end-to-end RAG latency for single queries on Desktop equipped with RTX4090, representing a typical local scenario. Figure 8 shows the latency reduction of TeleRAG across three datasets and two LLMs (Llama-3.2-3B and Llama-3-8B).
As shown in Figure 8, TeleRAG consistently outperforms the baseline across all tested configurations. With Llama-3.2-3B, TeleRAG achieves average speedups of $1.55\times$, $1.54\times$, and $1.49\times$ on NQ, HotpotQA, and TriviaQA, respectively. The best speedup of $2.11\times$ is achieved in the Iter-RetGen pipeline on HotpotQA, because this pipeline involves frequent retrieval operations and generally short LLM outputs, which enhances the relative impact of retrieval acceleration.
Another notable improvement is in the SubQuestion pipeline, where TeleRAG achieves approximately $1.85\times$ speedup across all datasets. This pipeline uses LLM-generated sub-questions and performs batched retrievals of 3 to 4 queries simultaneously. CPU-based retrieval suffers from limited parallelism in such scenarios, but TeleRAG efficiently utilizes GPU parallelism, significantly enhancing performance.
When deploying Llama-3-8B, speedups with TeleRAG slightly decrease compared to Llama-3.2-3B, primarily due to increased LLM latency and reduced available memory for prefetching. Still, TeleRAG achieves approximately $1.3\times$ speedup across datasets, with a peak improvement of $1.82\times$ for Iter-RetGen pipeline on HotpotQA. Achieving these results with only $3.75$ GB of remaining GPU memory (after accounting for Llama-3-8B (16 GB), the embedding model (1 GB), and other miscellaneous tensors) highlights TeleRAG’s robust capability in accelerating RAG inference even under tight GPU memory constraints.
Figure 10. Simulated throughput [^8] Results for 1 GPU are real, not simulated.of TeleRAG on NQ dataset with different number of H100 GPUs.
Multi-query throughput on H100. To evaluate the TeleRAG’s performance in batched inference, we evaluate the end-to-end throughput on Server (H100) in batch sizes 1, 2, 4, and 8. We show the results of six evaluated RAG pipelines with Llama-3-8B and Llama-2-13B in Figure 9.
[…]
Figure 11. Latency breakdown for Llama-3-8B on NQ with a H100 GPU in different batch sizes. nprobe is 256.
Figure 12. Comparison of end‑to‑end retrieval latency for two batching strategies: naive mini‑batching and similarity‑aware greedy grouping. The Greedy Group bars are too short to be visible.
Multi-GPU Throughput. We also evaluate TeleRAG in a multi-GPU system to further show the scalability. For the evaluation, we construct a simulation framework around our single-GPU implementation. We first define a global batch size that the whole system will process at once, and then we construct mini-batches to distribute among each GPU. To form a mini-batch, we perform a greedy search based on the embedding similarity of queries, grouping similar queries into the same mini-batch to maximize the cluster hit rate. We present the simulated throughput of TeleRAG with a global batch size of 128 and mini-batch size of 4 with 1 to 8 GPUs. As shown in Figure 8, TeleRAG’s throughput scales well with the number of GPUs. When compared to 1 GPU results, the averaged throughput enhancements of TeleRAG are $1.7\times$, $3.1\times$, and $5.4\times$ with $2$, 44, and $8$ GPUs, respectively.
5.4. Analysis and Sensitivity Study
Latency breakdown. We further show the latency breakdown of running RAG pipelines with Llama-3-8B and Llama-2-13B on a single H100 GPU in different batch sizes in Figure 11. From Figure 11, we can observe that LLM latency grows sub-linearly with larger batch sizes. However, the latency for Faiss retrieval on CPU grows linearly with the batch size, dominating the overall latency when batch size is large. These results echo our findings in Figure 9, and show limited scalability of CPU retrieval in serving scenario. In contrast, TeleRAG significantly accelerates across all batch sizes and achieves a higher speedup from $1.4\times$ to $1.8\times$ when the batch size increases from 1 to 8.
Ablation on mini-batch strategies. In our multi-GPU results, we implement a greedy grouping strategy for mini-batching based on queries’ similarity in a global batch. Here, we examine the benefits and overhead of this strategy in Figure 12, where we show the end-to-end retrieval latency for mini-batch size $4$ with four GPUs and a global batch of 128 queries. The grouping overhead is minimal and the overall retrieval latency of greedy grouping mini-batching consistently outperform naive mini-batching across all pipelines.
(a) Llama-3.2-3B with a RTX4090 GPU.
(b) Llama-3-8B with a H100 GPU.
Figure 13. Single-query retrieval speedup on NQ with different nprobe values.
Retrieval speedups across nprobe. Figure 13 shows the retrieval latency reduction on NQ. We observe consistent speedups for all nprobe values. The greatest speedups are achieved at nprobe 256, with average speedups of $7.21\times$ and $7.41\times$ on RTX4090 and H100, respectively. As a larger nprobe value is used, the retrieval performance of TeleRAG becomes constrained by missed clusters on the CPU, given our fixed prefetch budget across varying nprobe values. However, RAG inference with higher nprobe will result in longer latency spent on the retrieval and hence, TeleRAG still has a significant latency improvement against the CPU retrieval baseline.
Pipeline H100 (Llm3-8B) 4090 (Llm3-3B)
Budget Hit Rate Budget Hit Rate
HyDE 9 GB $91.95\%$ 7 GB $87.42\%$
SubQ 8 GB $79.04\%$ 7 GB $76.38\%$
Iter 5 GB $95.51\%$ 3 GB $84.59\%$
IRG 4 GB $59.52\%$ 2.5 GB $50.34\%$
FLARE 6 GB $79.35\%$ 3 GB $56.67\%$
S-RAG 3 GB $71.29\%$ 1.25 GB $29.96\%$
Table 4. The prefetch budget and corresponding averaged cluster hit rate for each pipeline and hardware setup on NQ dataset. The target retrieval nprobe is 256.
Prefetch budgets and cluster hit rates. Table 4 shows the prefetch budgets we set with the profile-guided approach on RTX4090 and H100 for NQ. It also presents the average cluster hit rate achieved with this prefetch budget. From the table, we can see that it generally achieves a high cluster hit rate (¿5050%) when TeleRAG has a large prefetching budget. For cases where the budget is less than 2 GB, we observe a relatively lower hit rate (¡5050%), limiting the benefits of reducing the CPU’s search workloads. However, as observed from Figure 8, TeleRAG achieves from $1.2\times$ to $1.6\times$ end-to-end speedups for these pipelines, thanks to the combined benefit of reducing CPU workload and utilizing the GPU to perform sorting on similarity distances.
6. Related Work
6.1. Systems for RAG
RAGCache (Jin et al., 2024a) proposes a caching system that stores KV caches from datastore chunks in an order-aware manner to improve the time to the first token. This approach only reduces prefill latency, leaving retrieval and decode times unaffected, despite the fact they typically dominate total latency (Agrawal et al., 2024). Moreover, it assumes repeated use of the same document across multiple requests, limiting its scalability over large data stores. Similarly, TurboRAG (Lu et al., 2024) precomputes KV caches from the data store, but it also only optimizes prefill latency. CacheBlend introduces a selective KV-cache fusion technique for RAG to reuse pre-computed caches (Yao et al., 2024).
RaLMSpec (Zhang et al., 2024b) proposes speculative retrieval and batched verification, Chameleon (Jiang et al., 2023d) proposes a CPU-GPU-FPGA heterogeneous architecture for accelerating the retrieval process, and PipeRAG (Jiang et al., 2024) is an algorithm-system co-design technique to overlap the retrieval and generation by modifying the algorithms. However, these works focus on the paradigm of RAG that retrieves documents once every few tokens and do not apply to modular RAG applications that are widely used now, as discussed in §2.1.
Speculative RAG (Wang et al., 2024b) introduces drafting by smaller LLMs to reduce RAG latency, but it does not target the retrieval latency. APIServe (InferCept) (Abhyankar et al., 2024) proposes a novel KV cache management strategy that can support the interception of LLM generation by other workloads, including retrieval. However, this work again does not focus on optimizing retrieval latency. EdgeRAG (Seemakhupt et al., 2024) reduces the memory requirement of retrieval by re-generating and caching document embeddings at runtime, targeting extremely resource-constrained environments. RAGO (Jiang et al., 2025) introduces an abstraction for RAG to automatically pick task placement, resource allocation, and batching policies.
Unlike all these prior works, we tackle system challenges of long retrieval latency and large memory requirement for modular RAG pipelines with large-scale datastores.
6.2. Systems for Compound LLM Applications
Apart from RAG, there is a growing interest in compound or agentic LLM applications, where multiple LLM calls and other applications are combined to serve complex functionalities (Zaharia et al., 2024; Berger and Zorn, 2024; Wang et al., 2024a). LLMCompiler (Kim et al., 2023a) is a framework that optimizes the execution of multiple functions in large language models by enabling parallel function calling. AI Metropolis (Xie et al., 2024) accelerates LLM-based multi-agent simulations with out-of-order execution. RAG is a specific type of application in this broader direction, and we propose systems techniques to optimize its execution latency, focusing on the characteristics of retrieval workload.
6.3. Vector Index
Vector index is a key component of RAG (Briggs et al., [n. d.]; Gao et al., 2023b), and many works have been proposed to improve their efficiency. ScaNN proposes an anisotropic quantization method for better efficiency (Guo et al., 2020). DiskANN and SPANN propose memory-disk hybrid indexing systems that work beyond the limitation of memory capacity (Chen et al., 2021; Jayaram Subramanya et al., 2019). Other prior works propose hardware acceleration of vector index with GPUs (Johnson et al., 2019), FPGAs (Zeng et al., 2022; Jiang et al., 2023a), TPU (Chern et al., 2022), or ray tracing hardware (Zhu, 2022; Liu et al., 2023). BANG proposes a method to scale graph-based ANN beyond GPU memory (Khan et al., 2024), and Rummy allows the index to scale beyond GPU memory capacity with reordered pipelining (Zhang et al., 2024a). These methods either require algorithm modifications or are bottlenecked by CPU-GPU bandwidth. Our proposal, focuses on the context of modular RAG applications, where queries for retrieval are usually generated by LLMs, and optimizes the latency and the GPU memory consumption without altering the algorithm of the IVF index.
7. Conclusion
In this paper, we introduced TeleRAG, an inference system that tackles the system challenges of RAG pipelins under latency-sensitive scenarios. By using lookahead retrieval, which overlaps data transfer and GPU computation for faster retrieval, a profile-guided approach to determine optimal prefetching data amount, and GPU-CPU cooperation, TeleRAG speeds up the end-to-end latency with minimal GPU memory requirement. Our evaluation shows that TeleRAG significantly improves performance compared to existing state-of-the-art solutions.